OpenStudy (anonymous):

What is the period of the function on y=4(sin3x)

OpenStudy (saifoo.khan):

Did you just wrote the answer as well?

OpenStudy (anonymous):

i messed up

OpenStudy (anonymous):

its a trap!

OpenStudy (saifoo.khan):

\[\frac{2 \pi}{n} = \frac{2 \pi}{3} = \checkmark\]